OpenStudy (anonymous):

Suppose a parabola has an axis of symmetry at x=-8, a maximum height of 2, and passes through the point (-7,-1). Write the equation of the parabola in vertex form

OpenStudy (anonymous):

Hello : )

OpenStudy (welshfella):

the vertex form of the parabola is y = a(x - b)^2 + c where x = b is the axis of symmetry, c is the max or minm value of y when x = b and a is some constant

OpenStudy (welshfella):

we know that axis of symmetry is x = -8 so that makes the equation y = -a(x -(-8)) + c which becomes y = a(x + 8^2 + c

OpenStudy (anonymous):

okay

OpenStudy (welshfella):

now you know that c = 2 and x = -7 when y = -1 so plug in these values and find a Thats it done

OpenStudy (anonymous):

okay, I got y=-3(x+8)^2+2

OpenStudy (welshfella):

thats a NOT -a v my mistake

OpenStudy (anonymous):

okay

OpenStudy (welshfella):

correct

OpenStudy (anonymous):

thank you, one more?

OpenStudy (welshfella):

plz re post

OpenStudy (anonymous):

k